A water pipe narrows from a radius of r1 = 4.99 cm to a radius of r2 = 2.75 cm. If the speed of the water in the wider part of the pipe is 1.31 m/s, what is the speed of the water in the narrower part? 


1
Expert's answer
2021-12-03T12:29:56-0500
"r_1^2v_1=r_2^2v_2\\4.99^2(1.31)=2.75^2v_2\\v_2=4.31\frac{m}{s}"